# Non-Gaussianities in primordial black hole formation and induced gravitational waves

### v1

@@ VI Conclusion and Discussion
In this paper we review the impact of the primordial non-Gaussinity in the curvature perturbation on the PBH formation as well as on the generation of scalar-induced GWs. We discussed some concrete inflation models to realize these interesting non-Gaussian curvature perturbations, and demonstrate how to calculate the PBH abundance by extended Press-Schechter formalism, as well as the leading order induced GWs. The main conclusion is that positive/negative non-Gaussianity will greatly enhance/suppress the PBH formation, while the induced GWs are only mildly enhanced for both cases. These two phenomena can be crosschecked in different mass windows and different frequency bands to probe the small-scale behavior of the primordial curvature perturbation, which are important scientific goals for many current and future experiments.
We focus on the simplest local-type non-Gaussianity, which gives $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ as a local function of Gaussian variable $\mathcal{R}_{\mathrm{g}}$. Various types of non-Gaussianities like equilateral, orthogonal, or folded shapes are also important, not only on the abundance but also on the clustering Shandera et al. (2013); Young et al. (2016); Tada and Yokoyama (2015); Young and Byrnes (2020); Suyama and Yokoyama (2019); Matsubara et al. (2019); Ragavendra et al. (2020); De Luca et al. (2021). Recently, Ref. Matsubara and Sasaki (2022) studied the bispectrum and trispectrum in the PBH formation systematically, and shows that in the narrow power spectrum case, the nonlinear parameter $f_{\mathrm{NL}}$ entering the calculation of PBH formation should be a linear combination of $f_{\mathrm{NL}}$ of local, equilateral, folded, and orthogonal shapes:
@@
which implies that other shapes of non-Gaussianity should also be taken into account when calculating the PBH abundance. Such non-Gaussianities are important for non-canonical kinetic terms, like for instance DBI inflation Silverstein and Tong (2004); Alishahiha et al. (2004), k-essence Armendariz-Picon et al. (1999); Garriga and Mukhanov (1999), or G-inflation Kobayashi et al. (2010).
@@
Our starting point is the $\delta N$ formalism, which assume that the Hubble patches we study are far enough from each other, such that each patch can be seen as a local Friedmann universe where the field value (and other perturbed quantities) have randomly distributed initial value among the vast number of such patches, and evolves in the later stage independently. This is called seperate universe approach. Such an approach works quite well in the slow-roll regime. However, in the beginning of ultra-slow-roll stage (marked by $\varphi_{\mathrm{i}}$), the gradient terms of both growing mode and decaying mode are rather important, which give rise to the famous $k^{4}$ growth Leach et al. (2001); Byrnes et al. (2019); Cole et al. (2022). Such a $k$-dependence cannot be recovered by seperate universe approach and the classical $\delta N$ formalism based on it Jackson et al. (2023). Fortunately, such a failure is only serious for $\varphi\lesssim\sqrt{\pi}\varphi_{\mathrm{i}}$ Domènech et al. (2023), which makes the seperate universe approach and the $\delta N$ formalsm still valid at the peak of the enhanced power spectrum, which is located at $\varphi_{\mathrm{p}}\approx\pi\varphi_{\mathrm{i}}$ Pi and Wang (2023). Apparently, if we extend the $\delta N$ formalism to include the spatial curvature on the initial slice taken slightly after horizon exit, the leading order correction of the gradient expansion can be taken into account Artigas et al. .
@@
We did not discuss the interesting case of negative non-Gaussianity, which is connected to Type II PBHs. Simple estimate shows that negative non-Gaussianity will significantly suppress the PBH abundance, which can be calculated in a perturbative manner Young and Byrnes (2013). For fully nonlinear curvature perturbation with a logarithmic dependence on the Gaussian variable, the PDF of $\mathcal{R}$ displays a Gumbel-like distribution (also called double exponential suppression):
$\mathbb{P}(\mathcal{R})\sim\exp\left(-c^{2}e^{2\lambda_{-}\mathcal{R}}\right),$ (86)
@@
where the $r\to 1$ limit of the curvaton scenario corresponds to $\lambda_{-}=3/2$. Direct calculation shows that it can also significantly suppress the PBH formation. This suppression was utilized to avoid the overproduction of sub-solar-mass PBHs when interpreting the nHz stochastoc GW background as the scalar-induced GWs Franciolini et al. (2023a); Liu et al. (2023). Recent study by peak theory indicates that the PBH formation for negative $f_{\mathrm{NL}}$ is quite difficult Kitajima et al. (2021), and numerical simulations show that there is a lower bound $f_{\mathrm{NL}}\gtrsim-1.2$ below which the Type I PBHs can not form, based on a monochromatic power spectrum Escrivà et al. (2022a). Around this regime the profiles of the density peaks become complicated, and the Type II PBH are crucial Uehara et al. (2024), which we did not touch here.
@@
In many models, the non-attractor behavior is excited by the sudden change of the slop of the potential, which itself is continuous. This actually means that the change of slope is realized in an energy scale much larger than $H$, i.e. $V^{\prime\prime}\gg H$. Another class of models display a discontinuity in the potential, which causes a sudden loss of the kinetic energy. Large non-Gaussianity can also be generated in such a case, which gives a highly asymmetric PDF of $\mathcal{R}$ Cai et al. (2022a, b); Kawaguchi et al. (2023). If the velocity is not enough, the inflaton may not overshoot the step in some Hubble patches. Recent study shows that such Hubble patches finally collapses to a PBH, as the energy density in the surrounding area become smaller than this patch and start to compress it. When the PBH from such bubbles are dominant, there is a large non-Gaussianity Escrivà et al. (2023).
@@
In principle, for non-quadratic potential, the equation of motion for the inflaton can not be solved analytically, and the analysis based on quadratic potentials in Section III.3 can not apply. However, inspired by the logarithmic relation of $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ and its exponential-taile PDF (61), it is rational to conjecture a more general PDF of $\mathcal{R}$ directly Nakama et al. (2016); Hooshangi et al. (2022, 2023). For instance, a simple guess is
@@
where $c$ is a positive constant. The PDF with $p<1$ is called heavy tailed, which is more effective in generating PBHs than the exponential-tail PDF. This extra efficiency is needed to generate enough supermassive PBHs to seed early galaxies, as the power spectrum of $\mathcal{R}$ on that scale is strongly constraints by $\mu$- and $y$-distortions, $\mathcal{P_{R}}\lesssim 10^{-5}$ Nakama et al. (2018); Carr and Silk (2018); Atal et al. (2020b); Liu and Bromm (2022); Gouttenoire et al. (2023). Such a heavy tail PDF can arise from self-interacting curvaton scenario Hooper et al. (2023), while a lighter tail of $p=3/2$ can be generated in the non-perturbative regime of inflation model with a ˙$\dot{\mathcal{R}}^{4}$ interaction Celoria et al. (2021).
Up to now, all the discussion above are based on the classical $\delta N$ formalism, which allows us to divide the field value $\varphi$ in a Hubble patch into “background” $\varphi_{0}$ and “perturbation” $\delta\varphi$. Such a split becomes difficult (if not completely impossible) when the perturbations are large and overwhelm the background value. As we commented, this happens when the inflaton is nearly static on a flat plateau due to the Hubble friction, while quantum diffusion should be taken into account Starobinsky (1986, 1982); Starobinsky and Yokoyama (1994). Although the exponential-tail PDF like (61) appears as the leading order of the PDF in the semi-classical limit, more complicated and bizarre PDF arises in the quantum regime Bullock and Primack (1997); Vennin and Starobinsky (2015); Pattison et al. (2017); Ezquiaga et al. (2020); Vennin (2020); Figueroa et al. (2021); Pattison et al. (2021); Figueroa et al. (2022); Animali and Vennin (2022); Ezquiaga et al. (2023). This is the topic of the next chapter Vennin and Wands (2024).
@@ Acknowledgement
I would like to thank Guillem Domènech, Cristiano Germani, Misao Sasaki, Teruaki Suyama, and Takahiro Tanaka for useful discussions, and to thank Ao Wang for helping me draw Fig 7. I also thank the hospitality of YITP, Kavli IPMU, IBS-CTPU, and KIAS during my visit when this paper was finalized, and thank all the participants of the 2nd International Workshop on Gravitational Waves and the Early Universe for the intriguing discussions. This work is supported in part by the National Key Research and Development Program of China Grant No. 2021YFC2203004, by Project No. 12047503 of the National Natural Science Foundation of China, by the JSPS Grant-in-Aid for Early-Career Scientists No. JP20K14461, by JSPS KAKENHI No. JP24K00624, and by the World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan.

### v2

@@ VI Conclusion and Discussion
In this paper we review the impact of the primordial non-Gaussinity in the curvature perturbation on the PBH formation as well as on the generation of scalar-induced GWs. We discussed some concrete inflation models to realize these interesting non-Gaussian curvature perturbations, and demonstrate how to calculate the PBH abundance as well as the induced GWs when non-Gaussianity is taken into account. Positive/negative non-Gaussianity can greatly enhance/suppress the PBH abundance, as they change the tail of the PDF significantly. However, the induced GW is mainly determined by the power spectrum, thus non-Gaussianity can not significantly change the induced GW spectrum. When fixing the PBH abundance, positive/negative non-Gaussianity only mildly suppresses/enhances the induced GWs. Therefore, the amplitude of the induded GW spectrum in the presence of abundant PBHs is a relatively robust prediction, which is an important scientific goal for many current and future experiments.
We only focus on the simplest local-type non-Gaussianity, which gives $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ as a function of Gaussian variable $\mathcal{R}_{\mathrm{g}}$. Scale dependence of such a relation is not considered here Passaglia et al. (2019); Ragavendra (2022); Özsoy and Tasinato (2022); Ragavendra and Sriramkumar (2023); Tasinato (2024). Especially, various types of non-Gaussianities like equilateral, orthogonal, or folded shapes are also important, not only on the abundance but also on the clustering Shandera et al. (2013); Young et al. (2016); Tada and Yokoyama (2015); Young and Byrnes (2020); Suyama and Yokoyama (2019); Matsubara et al. (2019); Ragavendra et al. (2020); De Luca et al. (2021). Recently, Ref. Matsubara and Sasaki (2022) studied the bispectrum and trispectrum in the PBH formation systematically, and shows that in the narrow power spectrum case, the nonlinear parameter $f_{\mathrm{NL}}$ entering the calculation of PBH formation should be a linear combination of $f_{\mathrm{NL}}$ of local, equilateral, folded, and orthogonal shapes:
@@
which implies that other shapes of non-Gaussianity should also be taken into account when calculating the PBH abundance. Such non-Gaussianities are important for inflation model with non-canonical kinetic terms, like for instance k-essence or G-inflation Lin et al. (2020); Yi et al. (2021); Zhang et al. (2021); Gao (2021); Choudhury et al. (2024).
@@
Our starting point is the $\delta N$ formalism, which assumes that the Hubble patches we study are far enough from each other, such that each patch can be seen as a local Friedmann universe where the field value (and other perturbed quantities) takes a random initial value and evolves independently in the later stage. This is called seperate universe approach, which works quite well in the slow-roll regime. However, in the beginning of ultra-slow-roll stage (marked by $\varphi_{\mathrm{i}}$), the gradient terms of both growing mode and decaying mode are rather important, which give rise to the well-known $k^{4}$ growth Leach et al. (2001); Byrnes et al. (2019); Cole et al. (2022). Such a $k$-dependence cannot be recovered by seperate universe approach and the classical $\delta N$ formalism based on it Jackson et al. (2023). Fortunately, such a failure is only serious for $\varphi\lesssim\sqrt{\pi}\varphi_{\mathrm{i}}$ Domènech et al. (2023), which makes the seperate universe approach and the $\delta N$ formalsm still valid at the peak of the enhanced power spectrum, which is located at $\varphi_{\mathrm{p}}\approx\pi\varphi_{\mathrm{i}}$ Pi and Wang (2023). Apparently, if we extend the $\delta N$ formalism to include the spatial curvature on the initial slice taken slightly after horizon exit, the leading order correction of the gradient expansion can be taken into account Artigas et al. .
@@
We did not discuss the interesting case of negative non-Gaussianity, which is connected to Type II PBHs. Simple estimate shows that negative non-Gaussianity will significantly suppress the PBH abundance, which can be calculated in a perturbative way Young and Byrnes (2013). For fully nonlinear curvature perturbation with a logarithmic dependence on the Gaussian variable, the PDF of $\mathcal{R}$ displays a Gumbel-like distribution Pi and Sasaki (2023b) (also called double exponential suppression Hooshangi et al. (2023)):
$\mathbb{P}(\mathcal{R})\sim\exp\left(-c^{2}e^{2\lambda_{-}\mathcal{R}}\right).$ (86)
@@
This can be realized, for instance, in the $r\to 1$ limit of the curvaton scenario which corresponds to $\lambda_{-}=3/2$. Direct calculation shows that it can significantly suppress the PBH formation. This suppression was utilized to avoid the overproduction of sub-solar-mass PBHs when interpreting the nHz stochastic GW background as the scalar-induced GWs Franciolini et al. (2023a); Liu et al. (2023); Choudhury et al. (2023a). Recent study by the theory of peaks indicates that the PBH formation for negative $f_{\mathrm{NL}}$ is quite difficult Kitajima et al. (2021), and numerical simulations show that there is a lower bound $f_{\mathrm{NL}}\gtrsim-1.2$ below which the Type I PBHs can not form, based on a monochromatic power spectrum Escrivà et al. (2022a). Around this regime the profiles of the density peaks become complicated, and the Type II PBH are crucial Uehara et al. (2024), which we did not touch here.
@@
In many models, the non-attractor behavior is excited by the sudden change of the slop of the potential, which itself is continuous. This actually means that the change of slope is realized in an energy scale much larger than $H$, i.e. $V^{\prime\prime}\gg H^{2}$ at $\varphi_{t}$. Another class of models display a discontinuity or a bump in the potential, which causes a loss of kinetic energy. Large non-Gaussianity can also be generated in such a case, which gives a highly asymmetric PDF of $\mathcal{R}$ Cai et al. (2022a, b); Briaud and Vennin (2023); Kawaguchi et al. (2023). If the initial velocity is not enough, the inflaton may not overshoot the step/bump in some Hubble patches. Recent study shows that such a Hubble patch finally collapse to a PBH, as the energy density in the surrounding area become smaller than this patch and starts to compress it. When $f_{\mathrm{NL}}\gtrsim 2.6$, the PBHs from such bubbles are dominant Escrivà et al. (2023); Caravano et al. (2024).
@@
In principle, for non-quadratic potential, the equation of motion for the inflaton cannot be solved analytically, and the analysis based on quadratic potentials in Section III.3 cannot apply. However, inspired by the logarithmic relation of $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ and its exponential-tail PDF (58), it is reasonable to conjecture a more general PDF of $\mathcal{R}$ directly Nakama et al. (2016); Hooshangi et al. (2022, 2023). For instance, a simple guess is
@@
where $c$ is a positive constant. The PDF with $10<p<$ 1 is called heavy tailed, which is more effective in generating PBHs than the exponential-tail PDF. This extra efficiency is needed to generate enough supermassive PBHs to seed early galaxies, as the power spectrum of $\mathcal{R}$ on that scale is strongly constrained by $\mu$- and $y$-distortions, $\mathcal{P_{R}}\lesssim 10^{-5}$ Kohri et al. (2014); Nakama et al. (2018); Carr and Silk (2018); Atal et al. (2020b); Liu and Bromm (2022); Gouttenoire et al. (2023). Such a heavy-tail PDF can arise from self-interacting curvaton scenario Hooper et al. (2023), while a lighter tail of $p=3/2$ can be generated in the non-perturbative regime of inflation model with a ˙$\dot{\mathcal{R}}^{4}$ interaction Celoria et al. (2021).
Up to now, all the discussions above are based on the classical $\delta N$ formalism, which allows us to divide the field value $\varphi$ in a Hubble patch into “background” $\varphi_{0}$ and “perturbation” $\delta\varphi$. Such a split becomes difficult (if not completely impossible) when the perturbations are large and overwhelm the background value. As we commented, this happens when the inflaton is nearly static on a flat plateau due to the Hubble friction, and quantum diffusion must be taken into account Starobinsky (1986, 1982); Starobinsky and Yokoyama (1994). Although the exponential-tail PDF like (58) appears as the leading order of the PDF in the semi-classical limit, more complicated and bizarre PDF arises in the quantum regime Bullock and Primack (1997); Vennin and Starobinsky (2015); Pattison et al. (2017); Ezquiaga et al. (2020); Vennin (2020); Figueroa et al. (2021); Pattison et al. (2021); Figueroa et al. (2022); Animali and Vennin (2022); Ezquiaga et al. (2023). This is the topic of the next chapter Vennin and Wands (2024).
@@ Acknowledgement
I would like to thank Diego Cruces, Guillem Domènech, Jaume Garriga, Cristiano Germani, Cristian Joana, Misao Sasaki, Teruaki Suyama, Takahiro Tanaka, Dong-Gang Wang, and Jianing Wang for their comments and discussions, and to thank Ryoto Inui and Ao Wang for helping me draw Fig 7. I also thank the hospitality of YITP, Kavli IPMU, IBS-CTPU, and KIAS during my visit when this paper was finalized, and thank all the participants of the 2nd International Workshop on Gravitational Waves and the Early Universe for the intriguing discussions. This work is supported in part by the National Key Research and Development Program of China Grant No. 2021YFC2203004, by Project No. 12047503 of the National Natural Science Foundation of China, by the JSPS Grant-in-Aid for Early-Career Scientists No. JP20K14461, by JSPS KAKENHI No. JP24K00624, and by the World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan.